SYMMETRY DEPENDENCE OF THE RO-VIBRONIC DISTRIBUTIONS OF THE ND_{2} A^{2}A_{1} FRAGMENTS FROM THE PHOTODISSOCIATION OF THE A STATES OF ND_{3} AND ND_{2}H AT 193.3 NM

\underline{G. DUXBURY}, Department of Physics, SUPA, John Anderson Building, University of Strathclyde, 107 Rottenrow, Glasgow G4 0NG, Scotland,UK; J.P. REID, School of Chemistry, University of Bristol, Bristol BS8 1TS.

A rotational and vibrational analysis has been made of the ND_{2} Ã^{2}A_{1}-X̃^{2}B_{1} emission spectrum produced from the ultraviolet laser induced dissociation of both jet cooled and room temperature deuterated ammonia, ND_{3}, and di-deutero ammonia, ND_{2}H. The pattern of the strong features in the emission spectra is very different in the fragmentation of ND_{3} and ND_{2}H, with a much wider range of angular momentum states being observed from the photolysis of the predissociative state of the unsymmetrical parent ND_{2}H. The analysis is based upon the earlier studies of the electronic spectrum of ND_{2}, and model calculation based upon the stretch-bender Renner-Teller Hamiltonian. The spectra consist of two types, transitions from a narrow distribution of high angular momentum states in the photolysis of ND_{3} and ND_{2}H, and in the photolysis of ND_{2}H strong emission from threshold states to three high energy regions of the X̃^{2}B_{1} state. The threshold states are in the third bending level, v2′=3, of the ND_{2} Ã^{2}A_{1} state, and have no angular momentum about the axis of least moment of inertia, Ka = 0.